# The DESI 2024 hint for dynamical dark energy is biased by low-redshift supernovae

### v3

@@ 2 Data and method
low-$z$ fluctuation debias $79.8\pm 7.7$ $-18.97^{+0.24}_{-0.19}$ $0.224^{+0.033}_{-0.056}$ $-0.64^{+0.24}_{-0.19}$ $-3.7^{+1.7}_{-2.0}$ $\sim 1.5\sigma$
Planck CMB + DESI BAO +
@@
DESI BAO $68.0^{+0.84}_{-0.71}$ $-19.406^{+0.024}_{-0.019}$ $0.308\pm 0.0076$ $-0.83^{+0.059}_{-0.070}$ $-0.70^{+0.30}_{-0.27}$ $\sim 2.5\sigma$
eBOSS BAO $67.3^{+0.61}_{-0.71}$ $-19.429^{+0.017}_{-0.020}$ $0.316\pm 0.0075$ $-0.88\pm 0.069$ $-0.45^{+0.29}_{-0.22}$ $\sim 1.7\sigma$
Planck CMB + DESI BAO + DESY5 debiased
@@
3 Analysis
3.1 Systematics identification
In Tab. 1, we present all constraints of the $w_{0}w_{a}$ model in this study, the first four groups of which are visualized in Fig. 1 for their marginalized $w_{0}-w_{a}$ distributions. In the left panel of Fig. 1, the base preference for dynamical DE from CMB+eBOSS BAO+PantheonPlus is $\sim 1.7\sigma$ (blue contour in the upper left), not significant yet. Either replacing the eBOSS BAO with DESI BAO (green contour in the upper right) or the PantheonPlus SNe with DESY5 SNe (purple contour in the bottom left) would increase preferences to $\sim 2.5\sigma$ and $\sim 2.9\sigma$, respectively, but still not decisive to claim dynamical DE. In particular, when getting rid of the low-$z$ SN sample, the significance drops from $\sim 2.9\sigma$ to insignificant $\sim 1.1\sigma$ (orange contour in the bottom left). Even though CMB+DESI BAO+DESY5 after both of the above replacements presents a more significant $\sim 3.5\sigma$ preference (red contour in the bottom right), getting rid of the low-$z$ SN sample would once again drop the significance from $\sim 3.5\sigma$ back to $\sim 1.7\sigma$ (orange dashed contour in the bottom right). On the contrary, both cases using low-$z$ SNe (blue dot-dashed contours in the bottom row) prefer a deviation from the $\Lambda$CDM model with a direction towards $w_{0}>-1$ and $w_{a}<0$. Therefore, a minor inconsistency emerges between low-$z$ SNe and DES-SN samples.
3.2 Systematics diagnosis
To diagnose systematics for low-$z$ SNe, we rebuild the magnitude-redshift relation
$m_{B,i}^{\mathrm{corr.}}$ $=5\lg\hat{d}_{L}(z_{i})-5a_{B},$ (1)
$-5a_{B}$ $\equiv M_{B}+5\lg\frac{c/H_{0}}{\mathrm{Mpc}}+25$ (2)
for both low-$z$ and DES-SN samples, where the dimensionless luminosity distance $\hat{d}_{L}(z_{i})=D_{L}(z)/(c/H_{0})=(1+z_{\mathrm{hel}})\int_{0}^{z_{\mathrm{HD}}}\mathrm{d}z^{\prime}/E(z^{\prime})$ is computed from the best-fit Planck $\Lambda$CDM fiducial model for the reduced Hubble parameter $E(z)^{2}\equiv H(z)^{2}/H_{0}^{2}=\Omega_{m}(1+z)^{3}+(1-\Omega_{m})$. Recall that all light-curve model corrections are used to standardize the apparent magnitude $m_{B,i}^{\mathrm{corr.}}$, and hence the intercept $-5a_{B}$ as a degeneracy parameter between $H_{0}$ and $M_{B}$ is expected to be a stable value for a single SNe compilation. The gist of $a_{B}$ diagnosis is to check whether the intercept $-5a_{B}$ (a systematics-sensitive and model-sensitive quantity) is consistent among different SNe samples. As any observational systematics in measuring $m_{B,i}^{\mathrm{corr.}}$ or theoretical systematics in modeling $5\lg\hat{d}_{L}(z_{i})$ induced by either uncounted peculiar velocity effect on the redshift $z_{i}$ measurement or even new physics model in $\hat{d}_{L}(z)$ would cause a deviation in the intercept $-5a_{B}=m_{B,i}^{\mathrm{corr.}}-5\lg\hat{d}_{L}(z_{i})$, we can quickly locate the redshift ranges of either systematics or new physics from redshift distributions of $-5a_{B}$.
Detailed $a_{B}$ diagnosis is visualized in Fig. 2. The high-quality DES-SN from a single telescope has the most stable $a_{B}$ distributions among its spanned redshift bins, perfectly agreeing with the Planck $\Lambda$CDM cosmology assumed in Fig. 2. By contrast, the multi-sample mixed low-$z$ SNe shows irregular fluctuations in $-5a_{B}$, and its weighted average is $\sim 0.043$ magnitude higher than that of DES-SN, which is qualitatively similar to the previous study by G. Efstathiou (2024) but with the $\sim 0.04$ mag. offset estimated on $\Delta m_{\mathrm{offset}}:=(m_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-m_{\mathrm{DESY5}}^{\mathrm{low}-z})-(m_{\mathrm{Pantheon+}}^{\mathrm{high}-z}-m_{\mathrm{DESY5}}^{\mathrm{high}-z})\approx(-0.05)-(-0.01)=-0.04$ mag. between PantheonPlus and DESY5 compilations in their low- and high-redshift ranges after assuming a constant $M_{B}$. As a response, M. Vincenzi et al. (2025) has explained partially this offset with improvements to the intrinsic scatter model and mass step estimate by reducing the offset from $\sim 0.04$ mag. to $\Delta\mu_{\mathrm{offset}}:=(\mu_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{low}-z})-(\mu_{\mathrm{Pantheon+}}^{\mathrm{high}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{high}-z})=(\mu_{\mathrm{Pantheon+}}^{\mathrm{low}-z}-\mu_{\mathrm{Pantheon+}}^{\mathrm{high}-z})-(\mu_{\mathrm{DESY5}}^{\mathrm{low}-z}-\mu_{\mathrm{DESY5}}^{\mathrm{high}-z})=0.019-0.043=-0.024$ mag., which can be further reduced to $-0.008$ mag. after considering a $\sim 0.016$ mag. difference in selection functions as elaborated by M. Vincenzi et al. (2025). However, such a negligible $-0.008$ mag. offset in $\Delta\mu_{\mathrm{offset}}$ is actually achieved between two different SN compilations, leaving an equivalent $0.019+0.008\sim 0.027$ mag. difference between low-$z$ and high-$z$ SNe within the single DESY5 compilation, which is still apparent in the difference of $-5a_{B}$ (after selection bias correction) between low-$z$ and high-$z$ parts of the single DESY5 compilation.
This $\sim 0.043$ magnitude discrepancy in the intercept $-5a_{B}$ within the single DESY5 compilation appears to emerge suddenly around $z\sim 0.1$ where the Universe is still homogeneous, hence we can safely rule out the local-scale inhomogeneous effects. This sudden transition also cannot originate from homogeneous-scale new physics since the PantheonPlus SNe, containing a larger number of low-$z$ well-calibrated SNe, does not detect any such low-$z$ transition. If a homogeneous-scale transition model could compensate for this intercept discrepancy within the DESY5 compilation, the same model would, in turn, reproduce a new intercept discrepancy within the PantheonPlus compilation. In addition, as revealed in the right panel of Fig. 2, the 1D probability density distribution (red dashed line) of low-$z$ $-5a_{B}$ is non-Gaussian, unlike the other two cases, suggesting the low-$z$ SNe suffers from large scatters and does not follow a consistent and tighten magnitude-redshift relation. As a comparison, the PantheonPlus SNe contains even more subsamples (18 samples) but still converges to a Gaussian distribution around the weighted average, indicating the well-corrected SNe have followed an accordant magnitude-redshift relation just as the DES-SN does. Therefore, the strange transition and large scatter in the low-$z$ intercept $-5a_{B}$ can only be attributed to systematics. Regardless of its origin, one can only use these low-$z$ SNe for cosmological analyses after fully de-biasing this mismatch intercept.
Figure 3: Left: The marginalized contours on parameters $w_{0}$ and $w_{a}$ for CMB+DESI BAO+DESY5 with different de-biasing corrections. Right: the same as the left panel but without CMB interference assuming $H_{0}=67.4$ km/s/Mpc.
3.3 Systematics corrections
The low-$z$ SNe systematics call for better sample selections and calibrations, which goes well beyond our current study. Based on the $a_{B}$ diagnosis, we propose a simple but rough correction procedure to the low-$z$ sample, which is, however, sufficient for our purpose to re-examine the preference for dynamic DE. Since the intercept $-5a_{B}$ measured from high-$z$ SNe is less affected by systematics from the peculiar velocity effect than that measured from low-$z$ SNe, the systematics correction for low-$z$ SNe is done by matching the low-$z$ intercept to the high-$z$ intercept. Specifically, the correction test includes three parts: (i) Smoothing the irregular fluctuations in the intercept $-5a_{B}$ for each redshift bin of low-$z$ SNe to match the weighted average of all low-$z$ SNe; (ii) Reducing the low-$z$ $m_{B}$ by $0.043$ magnitude so that the weighted average $-5a_{B}$ of low-$z$ SNe also matches the weighted average of DES-SN; (iii) Combining correction processes (i) and (ii) together to have a fully de-biased low-$z$ SN sample.
As a preliminary test, we find the $\sim 2.2\sigma$ preference of CMB+original low-$z$ SNe decreases to $\sim 1.5\sigma$ after roughly de-biasing the fluctuations in $-5a_{B}$ as shown in the first group of constraints in Tab. 1, which also slightly regularizes the $H_{0}-M_{B}$ contour (black unfilled contours) in the inset of the right panel of Fig. 1. Therefore, we add MCMC constraints of corrected low-$z$ and DESY5 in the last two groups of Tab. 1 and Fig. 3 with and without the interference from CMB. In the left panel of Fig. 3, correcting for the fluctuations in $-5a_{B}$ alone can cause little deviation in the $w_{0}-w_{a}$ contour, however, solely correcting for the low-$z$ discrepancy in $-5_{B}$ without debiasing the fluctuations first is already sufficient to reduce the preference for dynamical DE from $\sim 3.5\sigma$ to $\sim 1.6\sigma$, almost overlapping with the fully de-biased correction case with $\sim 1.5\sigma$. A larger reduction to $\sim 0.5\sigma$ and the same pattern also emerges in the right panel without CMB interference after assuming $H_{0}=67.4$ km/s/Mpc. A different $H_{0}$ prior only changes the degenerated constraints with respect to the sound horizon $r_{d}$ and absolute magnitude $M_{B}$ but still preserves the same reduced tension in the $w_{0}-w_{a}$ plane. Therefore, it is the $\sim 0.043$ magnitude discrepancy between low-$z$ and DES-SN that plays a leading role in driving the artificial preference for dynamical DE.
4 Conclusions and Discussions
Recent cosmological analyses from the DESI collaboration have attracted wide interest but also raised significant challenges and hot debate to the standard $\Lambda$CDM. In this letter, we locate the challenge from a systematics-dominated mirage with step-by-step data comparisons. The multi-sample mixed low-$z$ SNe in DESY5 suffering from large scatters fails to rebuild a consistent intercept of the magnitude-redshift relation and hence admits a $\sim 0.043$ magnitude discrepancy with respect to the DES-SN at higher redshifts, which drives the spurious dynamical behavior of DE. We roughly correct this discrepancy in the intercept for the low-$z$ SN samples and find the corrected datasets favor dynamical DE only at $0.5\sigma\sim 1.5\sigma$ level, not significant at all to call for dynamical DE. Therefore, the success of the standard $\Lambda$CDM has not been overturned.
Although we have identified the inclusion of low-$z$ SNe samples is largely responsible for the immature claim of dynamical DE, the DESI BAO data alone indeed shows some deviations from the Planck-$\Lambda$CDM model at higher redshift especially the luminous red galaxies (LRG) between $0.80.6<z<$ 0.8 (A. G. Adame et al., 2024a; M. Abdul Karim et al., 2025; U. Andrade et al., 2025), which have also been considered recently by Z. Wang et al. (2024); G. Liu et al. (2024) as the possible drives for the preference of dynamical DE. Nevertheless, we believe this is an independent indication for dynamical DE (K. Lodha et al., 2025), different from current study when including SNe Ia, and whether such a quintom-like DE (B. Feng et al., 2006) should be replaced by modified gravity (G. Ye et al., 2024) or thawing DE (S.-F. Chen et al., 2024; W. J. Wolf et al., 2025, 2024) or even both (e.g., R.-G. Cai et al., 2021) is highly worth further pursuing.
We thank Tamara Davis, George Efstathiou, Bin Hu, Yipeng Jing, Cheng Zhao, Gong-Bo Zhao, Jian-Qi Liu, and Yan-Hong Yao for helpful discussions. This work is supported by the National Key Research and Development Program of China Grants No. 2021YFA0718304, No. 2021YFC2203004, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 11821505, No. 12235019, No. 11991052, No. 12047503, No. 12073088, and No. 11947302, the Science Research Grants from the China Manned Space Project with No. CMS-CSST-2021-B01 (supported by China Manned Space Program through its Space Application System), and the Postdoctoral Fellowship Program of CPSF. We also acknowledge the use of the HPC Cluster of ITP-CAS. {contribution} All authors contributed equally to the manuscript.

### v4

@@ 4 Conclusions and Discussions
Planck CMB + DESI BAO DR1 +
@@
DESI BAO DR1 $68.0^{+0.84}_{-0.71}$ $-19.406^{+0.024}_{-0.019}$ $0.308\pm 0.0076$ $-0.83^{+0.059}_{-0.070}$ $-0.70^{+0.30}_{-0.27}$ $\sim 2.5\sigma$
eBOSS BAO DR16 $67.3^{+0.61}_{-0.71}$ $-19.429^{+0.017}_{-0.020}$ $0.316\pm 0.0075$ $-0.88\pm 0.069$ $-0.45^{+0.29}_{-0.22}$ $\sim 1.7\sigma$
Table 2: Constraints on $w_{0}w_{a}$ model with and without DESI Y1/Y3 BAO before and after rebuilding $a_{B}$ consistency.
Planck CMB
Datasets $H_{0}$ $M_{B}$ $\Omega_{m}$ $w_{0}$ $w_{a}$ Tensions
low-$z$ fluctuation debiased $79.8\pm 7.7$ $-18.97^{+0.24}_{-0.19}$ $0.224^{+0.033}_{-0.056}$ $-0.64^{+0.24}_{-0.19}$ $-3.7^{+1.7}_{-2.0}$ $\sim 1.5\sigma$
Planck CMB + DESI Y1 BAO
Datasets $H_{0}$ $M_{B}$ $\Omega_{m}$ $w_{0}$ $w_{a}$ Tensions
DESY5 $67.25\pm 0.72$ $-19.370\pm 0.017$ $0.316\pm 0.0077$ $-0.73\pm 0.069$ $-1.01\pm 0.32$ $\sim 3.5\sigma$
@@
Planck CMB + DESI Y3 BAO +
Datasets $H_{0}$ $M_{B}$ $\Omega_{m}$ $w_{0}$ $w_{a}$ Tensions
DESY5 $66.97\pm 0.66$ $-19.379\pm 0.017$ $0.318\pm 0.0069$ $-0.76\pm 0.06$ $-0.86\pm 0.27$ $\sim 3.7\sigma$
low-$z$ fluctuations debiased $66.76^{+0.74}_{-0.66}$ $-19.382\pm 0.017$ $0.320^{+0.0066}_{-0.0075}$ $-0.74\pm 0.068$ $-0.88^{+0.32}_{-0.24}$ $\sim 3.5\sigma$
low-$z$ discrepancy debiased $68.30\pm 0.69$ $-19.379\pm 0.017$ $0.305\pm 0.0067$ $-0.90^{+0.057}_{-0.065}$ $-0.45^{+0.30}_{-0.23}$ $\sim 1.6\sigma$
low-$z$ fully debiased $68.39\pm 0.78$ $-19.377\pm 0.017$ $0.305^{+0.0061}_{-0.0073}$ $-0.91\pm 0.063$ $-0.43^{+0.28}_{-0.23}$ $\sim 1.6\sigma$
DESI Y3 BAO +
Datasets $r_{d}$ $M_{B}$ $\Omega_{m}$ $w_{0}$ $w_{a}$ Tensions
DESY5 $146.1\pm 1.6$ $-19.368\pm 0.014$ $0.317^{+0.039}_{-0.018}$ $-0.787^{+0.074}_{-0.088}$ $-0.69\pm 0.69$ $\sim 2.0\sigma$
low-$z$ fluctuations debiased $146.1\pm 1.6$ $-19.369\pm 0.014$ $0.308^{+0.047}_{-0.013}$ $-0.788^{+0.075}_{-0.087}$ $-0.60^{+0.65}_{-0.78}$ $\sim 1.9\sigma$
low-$z$ discrepancy debiased $149.6\pm 1.6$ $-19.418\pm 0.013$ $0.238^{+0.098}_{-0.024}$ $-0.943^{+0.069}_{-0.11}$ $0.46^{+0.66}_{-0.27}$ $\sim 0.8\sigma$
low-$z$ fully debiased $149.7\pm 1.6$ $-19.419\pm 0.013$ $0.248^{+0.086}_{-0.023}$ $-0.954^{+0.069}_{-0.099}$ $0.42^{+0.69}_{-0.29}$ $\sim 0.7\sigma$